OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence.a16 = 21, a17 = -1

OpenStudy (anonymous):

please help! these are the answers I'm given: an = 351 + 22(n + 1) an = 351 - 22(n + 1) an = 351 + 22(n - 1) an = 351 - 22(n - 1)

OpenStudy (blank ):

an = 351 - 22(n - 1)

OpenStudy (anonymous):

thank you!

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)